OpenStudy (anonymous):

susan is making brownies. she uses 4 cups of flour to make 24 brownies. She wants to make 33 brownies, how many cups of flour does she need?

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

for 24 brownies 4 cups flour required for 1 brownie 4/24 cups required so for 33 brownies 4/24 x 33 ie 132/24 = 5.5 cups required

OpenStudy (anonymous):

I had replied earlier also but u apparently did not see it

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

pls click the Good Answer button to show status of question as solved

OpenStudy (anonymous):

A recipe for 2.5 cups of flour will make 20 cupcakes. How many cups of flour are needed to make 50 cupcakes?

OpenStudy (anonymous):

6.25 cups

OpenStudy (anonymous):

Thank you again!

OpenStudy (anonymous):

Using same logic, for 20 cupcakes u need 2.5 cups for 1 cupcake u need 2.5/20 cups for 50 cupcakes u need 2.5/20 x 50 = 125/20 = 6.25 cups